OpenStudy (anonymous):

The area of this cylinder's base is 150 square inches. The height of the cylinder is 6 inches. What is the volume of the cone that has the same base area and height? Show the steps of your solution and explain your work.

OpenStudy (ghazi):

pi * r^2 = 150 r^2= 150/3.14=47.74 \[r=\sqrt{47.74}\]=6.911

OpenStudy (ghazi):

volume of cone = 1/3*pi*r^2*height= 1/3* 3.14*47.74*6=296.80